Company AG produces and sells rocket parts. The company presents the following data related to its variable and fixed indirect costs. Both costs are allocated using machine hours of Laser cutter.

Budgeted:

-Variable Indirect Cost Rate $ 200.00 per hour

-Indirect Fixed Cost Rate $ 240.00 per hour

-Hours required per piece of rocket 1.5 hours

-Units to produce and sell for 2018 5,000 pieces

-Total Indirect Fixed Cost $ 1,800,000

Results for the year 2018:

-Units produced and sold 4,800 pieces

-Hours of used laser cutter 8,400 hours

-Total variable indirect costs $ 1,478,400

-Total fixed indirect costs $ 1,832,200

Find the following variations related to the indirect cost:

a.Spending Variance the variable part

b. Efficiency Variance

c. Spending Variance the fixed part

d. Production volume Variance
Answer #1

Solution a:

Actual rate of variable indirect cost = $1,478,400 / 8400 = $176 per hour

Spending variance of variable part = (SR - AR) * AH = ($200 - $176) * 8400 = $201,600 Favorable

solution b:

Efficiency variance - Variable part = (SH - AH) * SR = (4800*1.5 - 8400) * $200 = $240,000 Unfavorable

Solution c:

Spending variance - fixed part = budgeted fixed indirect cost - Actual fixed indirect cost = $1,800,000 - $1,832,200 = $32,200 Unfavorable

Solution d:

Production volume variance = Fixed overhead applied - Budgeted fixed overhead

= (4800*1.5*$240) - $1,800,000 = $72,000 Unfavorable